# Atomically clean free-standing two-dimensional materials through heating in ultra-high vacuum Supporting Information for Atomically clean free-standing two-dimensional materials through heating in ultra-high vacuum

### v1

@@ Abstract
Philipp Irschik1,∗, David Lamprecht1,2, Shrirang Chokappa1, Clemens Mangler1, Carsten Speckmann1, Thuy An Bui1, Manuel Längle1, Lado Filipovic2, Jani Kotakoski1,∗
1Faculty of Physics, University of Vienna, Boltzmanngasse 5,
1090 Vienna, Austria
2Institute for Microelectronics, TU Wien, Gußhausstraße 25-29, 1040 Vienna, Austria
@@ Introduction
Two-dimensional (2D)materials, such as monolayer graphene and hexagonal boron nitride (h-BN)possess a much larger surface-to-volume ratio compared to bulk materials, leading to an amplified effect of the often undesired surface contamination, which can result in the degradation of a variety of the materials’ intrinsic properties, including their charge carrier mobility,[1, 2, 3] wettability,[4] and thermal conductivity.[5, 6] Additionally, the extent of its impact can vary between samples or even within different areas of the same specimen due to the lack of control over it, hindering reproducibility in device operations.[7] Furthermore, even a thin film of adsorbed contaminants can obstruct the underlying material, with clean areas often being only a few to tens of nanometers in size.[8] This can be detrimental to contrast-based imaging techniques, such as (scanning) transmission electron microscopy, particularly for atomically thin free-standing specimens, where the signal of interest comes only from one to a few layers of atoms, and clean areas larger than a few nanometers are often desired.[9] These are necessary, as the observed properties may actually stem from interactions with contamination[10, 11, 12] rather than the material itself,[13, 14] and could lead to misinterpretation and misreporting of the properties of the studied materials.
@@
Even though in-situ methods for sample preparation of 2D materials have been developed,[15] it is still common to use ex-situ procedures, such as mechanical exfoliation, molecular beam epitaxy, or chemical vapor deposition (CVD), where the prepared samples are exposed to ambient conditions, either during or after the synthesis process,[16, 17, 18] resulting in exposure to airborne hydrocarbon contamination, silicon, water vapor, and molecular oxygen already after a few minutes.[4, 19, 20, 21, 22] Also, transferring the materials onto other surfaces is typically needed for further device fabrication or applications, a process that often involves direct contact with polymers like poly(methyl methacrylate) (PMMA).[23, 24] While organic solvents like acetone and chloroform[25] can be used to dissolve PMMA after the transfer, they often fail to remove it completely, and can thus leave a thin film of polymer adsorbed on the surfaces,[26, 27, 28] obstructing the underlying materials, and degrading their properties.[3]
@@
This ubiquitous presence of surface contamination has incentivized the development of various post-preparation cleaning methods. While specialized cleaning techniques like cleaning in activated carbon,[19, 29] electrolytic cleaning,[30] or annealing with the aid of metal catalysts[31, 32] have been explored, they are not suitable for all applications and often require specialized device fabrication. Thus, methods such as mechanical cleaning, plasma cleaning, laser cleaning, current-induced annealing, and thermal annealing are still often used.[33].
@@
While mechanical cleaning methods can produce atomically clean areas by breaking the van der Waals bonds between the contamination and the underlying material using tips commonly used in atomic force microscopy (AFM),[34, 22] this is an inherently slow process and thus does not scale well to larger areas. Plasma cleaning can be effective in etching surface-bound polymer contamination,[28] however, a successful etching process relies on carefully chosen parameters for the plasma energy and density, as well as treatment duration, to not damage the underlying material in the process.[35, 36] Cleaning techniques using laser ablation can also effectively remove contamination, but are limited to the laser spot size, restricting their application on the larger scale,[37, 38] and methods like current-induced annealing are limited to electrically conductive materials and require electrode attachment.[39, 40, 41]
Therefore, thermal solutions for sample cleaning are often used due to their rather simple installation and ease of use. Many setups also allow annealing in controlled ambient environments, such as inert, reducing, and oxidizing atmospheres, as well as vacuum, the selection of which can have a drastic effect not only on the cleaning process but also on the integrity of the 2D material, with electrical doping,[42, 43] defect formation,[44, 45] induced strain,[46, 47] and graphitization of carbonaceous contamination[48] being common side effects of thermal annealing processes.
@@
For graphene, even at temperatures as high as 500 °C, inert atmospheres have been reported to be ineffective at completely removing PMMA residue,[49] and can turn it into covalently bonded amorphous carbon, which makes it difficult to remove.[50, 30] Reducing atmospheres can facilitate depolymerization of polymer residue when annealed at 400 °C or higher, with Ar/H2 being particularly effective in removing residual conjugated carbon systems by hydrogenation of the C=C bonds.[42] However, it tends to induce rehybridization of graphene from sp2 to sp3, resulting in amorphous carbon,[44, 46] as well as strong p-type doping.[43] Annealing in oxidizing atmospheres, while being effective in decomposing amorphous carbon contamination,[51] can cause the formation of cracks in graphene at temperatures as low as 200 °C,[44] and still fails to remove all contaminants even at 500 °C.[38] Also, it has been found that the oxidative strength of the atmosphere has a strong effect on the etching characteristics, as NO2 seems to completely etch graphene away when annealed above 500 °C, while CO2 preserves it[45]. Thermal annealing in vacuum, while reported to be unable to remove contamination completely,[2, 52, 27, 49] seems to introduce the least amount of lattice defects compared to other thermal annealing treatments due to the lack of reactive species in the annealing atmosphere.[25] In addition, the amount of induced hole doping and strain is reported to be minimal,[43, 47] while providing satisfactory cleaning results.[49, 38] The purity of the vacuum must be kept in mind though, as annealing under better vacuum conditions reportedly yields cleaner specimens.[50] Pressure values in the range have been shown to be ineffective at reliably cleaning graphene surfaces,[49] compared to annealing in high vacuum[25] or ultra-high vacuum (UHV),[2, 52, 27, 47] where large atomically clean surfaces have been achieved,[38] making UHV annealing a suitable candidate for contamination removal without harming the integrity of the underlying material.
@@
Despite this established understanding of polymer degradation as well as hydrocarbon species desorption[53, 54] on graphene surfaces, there still is a lack of reporting on the cleaning of monolayer h-BN despite its various projected applications, such as being a dielectric for microelectronic devices,[55, 56, 57, 58, 59] and for quantum information and sensor technology due to its quantum light emission across a wide spectral range.[60, 61, 62, 63, 64]. Particularly, the latter relies on atomically clean surfaces to reliably conduct atomically-precise defect creation[65, 66, 67] or carbon doping[68, 69, 70] to create active sites for single photon emission. While surface contamination reduction via thermal annealing of h-BN has previously been utilized, including annealing in oxidizing environments,[71, 72, 73, 74] and UHV,[75, 76] the reports either lack information on the structure and abundance of contaminants as a function of annealing temperature, or do not employ atomically resolved characterization techniques, leaving open the possibility of a thin layer of contaminants remaining on the surface.
@@
In this work, we investigate the effect of thermal annealing in UHV on CVD-grown and wet-transfer-assisted free-standing monolayer graphene and h-BN, using a custom-built UHV heating chamber that is connected to a scanning transmission electron microscope through a vacuum transfer line. This ensures that the cleaning effect can be directly characterized without introducing airborne contaminants in the process. Contrary to commonly used spectroscopic assessment methods with limited spatial resolutions, including Raman spectroscopy,[43, 30] X-ray photoelectron spectroscopy,[2, 42, 49, 35] and time-of-flight secondary ion mass spectrometry,[27, 47, 77] we employ atomically resolved scanning transmission electron microscopy (STEM)to identify the presence of contamination on monolayer graphene and h-BN surfaces as a function of temperature down to the atomic level. Using annular dark field (ADF)-STEM imaging, we can identify the position, shape, and abundance of contamination due to its higher scattering intensity compared to atomically clean free-standing monolayer regions. Using contrast-based image thresholding and segmentation to separate contaminated from clean areas, we determine the relative amount of surface area covered by contamination, a procedure that is carried out for different transfer methods and baking temperatures, correlating the relative clean monolayer area of graphene and h-BN with annealing temperature. This is complemented with electron energy loss spectroscopy (EELS)measurements before and after the annealing steps, giving insight into the composition of the contamination that survives the heating process, which is one of the main limitations for obtaining truly residue-free materials.
@@ Methods
After investigating the sample quality and noticing the large amount of metallic residue on all samples transferred this way, both before and after the annealing steps, further graphene samples were prepared via electrochemical delamination[78] using commercial monolayer graphene on Cu with PMMA coating (Graphenea, Inc.). To delaminate the samples, after cutting them to size as described above, a platinum anode was placed into NaOH (SigmaAldrich, ACS reagent, , pellets) aqueous solution (1 m), with the Cu/graphene/PMMA stack on Inox 02 tweezers acting as the cathode. All samples were delaminated at a voltage of 4 V. Afterwards, the delaminated graphene/PMMA stack was washed in DI-water to remove residual NaOH contamination, after which it was fished out with a TEM grid with a Quantifoil holey amorphous carbon membrane (Quantifoil R 1.2/1.3 Au grid), followed by a hot acetone bath for 1 h, with a final wash in IPA for 45 min. One additional sample was transferred onto a holey silicon nitride support film (PELCO, diameter, silicon support structure, 0.5×0.5 mm window, 200 nm membrane thickness, pore size) using the same transferring steps.
Hexagonal boron nitride— After sourcing monolayer h-BN from three different suppliers, we chose monolayer h-BN on copper from SigmaAldrich, cut into small squares, which was spin-coated with PMMA (ALLRESIST AR-P 642.04, 200K) for 15 s at 500 RPM, followed by 15 s at 1000 RPM, and finally 60 s at 2000 RPM, at a ramp of 100 RPM/s. The Cu/h-BN/PMMA stack was put on a glass slide, where the PMMA was left to cure on a hot plate at 150 °C for 30 min in air. After the PMMA had hardened, less than 1 mm was cut off all edges of the square to remove PMMA spillover to the other side of the foil. The subsequent delamination and washing steps were analogous to the graphene samples transferred via electrochemical delamination as described above.
@@
All annealing experiments were carried out in a custom-built UHV heating chamber integrated into the CANVAS system at the University of Vienna,[93] with a base pressure typically around or lower. The samples are annealed via resistive heating of a tungsten filament that is placed out of line of sight of the samples to avoid locally high temperatures and deposition of material coming off the wire. A more thorough description, including design sketches and used materials, can be found in the Supporting Information.
@@
We use commercial CVD-grown monolayer graphene and h-BN, which were transferred onto TEM grids using a PMMA-assisted wet transfer process (see Methods). The majority of the PMMA residue was removed from the surface by placing the samples into an acetone bath, followed by a further washing step in IPA. After preparation, the samples were inserted into the CANVAS system at the University of Vienna,[93] which allows transporting the samples between a customized aberration-corrected scanning transmission electron microscope (Nion UltraSTEM 100),[80] and a custom-built UHV heating chamber without breaking the near-UHV conditions in between. This is integral to assessing the samples’ cleanliness since this allows the cleaning effect to be directly characterized without intermediary air exposure. Upon insertion into the loadlock of the CANVAS system, they were subjected to a routine bake at 160 °C for over 10 h in high vacuum to remove water and reduce the amount of surface contamination to avoid contaminating the UHV system. Using the Nion UltraSTEM 100, all samples were investigated for coverage and cleanliness before the UHV annealing step, after which they were transferred to the UHV heating chamber and annealed at different temperatures for 3 h. After the baking procedure had finished, the samples were immediately returned to the microscope for post-heating characterization to minimize the effect of electron-beam-induced hydrocarbon deposition.[81]
To assess the cleanliness of the samples, both before and after the thermal treatment, up to ten typical holes ( nominal diameter) in the sample support were imaged. The holes were chosen based on average and characteristic features resulting from each heating process. Different holes in the sample support were selected before and after éach heating step, as the dissociation and subsequent cross-linking of the hydrocarbon contaminants from prior e-beam exposure interferes with heating-induced contamination removal.[82] Approximately 20 arbitrarily chosen, non-overlapping areas with nominal FOV of 128×128 nm2 were captured within each hole, acquiring a total of ca. 150 - 200 frames per temperature and sample. Each time we moved to the position of the subsequent frame, the electron beam was re-focused onto the sample surface to obtain lattice resolution, minimizing out-of-focus features in the contamination at the interface to the material’s surface. We note that no electron-beam-induced hydrocarbon deposition has been observed during any measurement, and refer to the Supporting Information for a more thorough discussion.
@@
All as-prepared Easy Transfer Graphene samples exhibited a high amount of metallic contamination in the form of nanosized iron and copper clusters in addition to carbonaceous contamination covering practically the entire surface. While atomically clean monolayer regions could be found, they were small in size, typically less than 10 nm in diameter, and accounted for less than 5% of the free-standing regions (Figure 1a, blue dots). Example areas of contaminated and cleaned Easy Transfer Graphene are shown in Figure 1b. Heating to 250 °C did result in some removal of organic residue, increasing the average exposed clean area to 35%. The sizes of the now uncovered clean areas remained small (around 20 - 30 nm in diameter), indicating that higher temperatures are needed. Heating to 320 °C uncovered more of the surface area, however, we identified thin hydrocarbon contamination networks all over the surface with only a marginal increase of the uncovered clean area compared to the 250 °C treatment. While annealing at 450 °C removed the majority of the carbonaceous contamination, resulting in a 70% relative clean monolayer area with the widths of clean areas exceeding being easily found, we identified networks of metal clusters embedded in carbon-rich contamination as the primary residual contamination. Raising the annealing temperature to 550 °C did not increase the unveiled clean area, which remained at around 70%. This not only indicates that the achievable cleanliness via UHV annealing is heavily limited by the residual metal contaminants introduced during the sample growth and transfer process, but also that temperatures above 400 °C appear to be enough to remove almost all of the carbonaceous contamination that is not directly bound to metal clusters.
@@
The remaining residues after annealing at 450 °C consist mainly of copper, but also iron nanoclusters (Supplementary Figure 12, Supporting Information), embedded in carbon-rich contamination, primarily located at graphene grain boundaries, shown in Figure 2a. These metal clusters also form aggregates with a higher density of metallic residue, which are not necessarily attached to grain boundaries as seen in Figure 2b, from which dendritic carbon-rich structures appear to emerge. Supplementary Figure 13 (Supporting Information) shows that the metal contaminants are rich in Fe and Cu. We also identify nanosized holes in the material with typical sizes between 15 and 20 nm (Figure 2c). However, based on the relatively low abundance of these holes, we presume they were already present in the material before annealing and were covered by contamination removed during the heating process, instead of being annealing-induced damage. Atomically clean areas that would otherwise be thousands of nm2 in size contain some isolated small carbon contamination patches (Figure 2d), likely attached to point defects in the material. Contaminants trapped between layers that folded over on each other during the sample transfer underwent minimal cleaning, if any at all, as they cannot penetrate and desorb from the graphene layers they are sandwiched between (Figure 2e) and can only undergo diffusion processes.
@@
Upon investigating the pre-cleaned samples, we observed that the quality of the transferred h-BN was inferior to that of graphene made via the same transfer method. Firstly, the suspended regions showed more holes and folds compared to graphene, likely originating from the transfer process. Secondly, a large fraction of the surface was covered in stacked multilayer structures, most of which appeared to emerge from SiO2 nanoclusters (Supplementary Figure 14, Supporting Information). Given that almost all of these multilayer structures were found around said clusters, we presume that they were introduced during the sample growth, possibly from evaporation of SiO2 from the quartz tube during the CVD growth process,[83] rather than during storage or transfer. Together, these two sources contributed to an estimated monolayer coverage of only around 50% (Supplementary Figure 14, Supporting Information), the highest among the three h-BN suppliers considered for this study (Supplementary Figure 15, Supporting Information), compared to practically 100% for graphene, despite the similar sample preparation procedures. This is also reflected in the strong variation of measured clean monolayer area (Figure 3a) across all temperatures, which is largely due to differences in monolayer coverage per imaged area, rather than local variations in residual contamination at a given temperature.

### v2

@@ Abstract
Philipp Irschik1,2,∗, David Lamprecht1,3, Shrirang Chokappa1,2, Clemens Mangler1, Carsten Speckmann1, Thuy An Bui1,2, Manuel Längle1, Lado Filipovic3, Jani Kotakoski1,∗
1University of Vienna, Faculty of Physics, Boltzmanngasse 5, 1090 Vienna, Austria
2University of Vienna, Vienna Doctoral School in Physics, Boltzmanngasse 5, 1090 Vienna, Austria
3Institute for Microelectronics, TU Wien, Gußhausstraße 27-29/E360, 1040 Vienna, Austria
@@ Introduction
Two-dimensional (2D)materials, such as monolayer graphene and hexagonal boron nitride (h-BN)possess a much larger surface-to-volume ratio compared to bulk materials, leading to an amplified effect of the often undesired surface contamination, which can result in the degradation of a variety of the materials’ intrinsic properties, including their charge carrier mobility[1, 2, 3], wettability[4], and thermal conductivity[5, 6]. Additionally, the extent of its impact can vary between samples or even within different areas of the same specimen due to the lack of control over it, hindering reproducibility in device operations[7]. Furthermore, even a thin film of adsorbed contaminants can obstruct the underlying material, with clean areas often being only a few to tens of nanometers in size[8]. This can be detrimental to contrast-based imaging techniques, such as (scanning) transmission electron microscopy, particularly for atomically thin free-standing specimens, where the signal of interest comes only from one to a few layers of atoms, and clean areas larger than a few nanometers are often desired[9]. These are necessary, as the observed properties may actually stem from interactions with contamination[10, 11, 12] rather than the material itself[13, 14], and could lead to misinterpretation and misreporting of the properties of the studied materials.
@@
Even though in-situ methods for sample preparation of 2D materials have been developed[15], it is still common to use ex-situ procedures, such as mechanical exfoliation, molecular beam epitaxy, or chemical vapor deposition (CVD), where the prepared samples are exposed to ambient conditions, either during or after the synthesis process[16, 17, 18], resulting in exposure to airborne hydrocarbon contamination, silicon, water vapor, and molecular oxygen already after a few minutes[4, 19, 20, 21, 22]. Also, transferring the materials onto other surfaces is typically needed for further device fabrication or applications, a process that often involves direct contact with polymers like poly(methyl methacrylate) (PMMA)[23, 24]. While organic solvents like acetone and chloroform[25] can be used to dissolve PMMA after the transfer, they often fail to remove it completely, and can thus leave a thin film of polymer adsorbed on the surfaces[26, 27, 28], obstructing the underlying materials, and degrading their properties[3].
@@
This ubiquitous presence of surface contamination has incentivized the development of various post-preparation cleaning methods. While specialized cleaning techniques like cleaning in activated carbon[19, 29], electrolytic cleaning[30], or annealing with the aid of metal catalysts[31, 32] have been explored, they are not suitable for all applications and often require specialized device fabrication. Thus, methods such as mechanical cleaning, plasma cleaning, laser cleaning, current-induced annealing, and thermal annealing are still often used[33].
@@
While mechanical cleaning methods can produce atomically clean areas by breaking the van der Waals bonds between the contamination and the underlying material using tips commonly used in atomic force microscopy (AFM)[34, 22], this is an inherently slow process and thus does not scale well to larger areas. Plasma cleaning can be effective in etching surface-bound polymer contamination[28], however, a successful etching process relies on carefully chosen parameters for the plasma energy and density, as well as treatment duration, to not damage the underlying material in the process[35, 36]. Cleaning techniques using laser ablation can also effectively remove contamination, but are limited to the laser spot size, restricting their application on the larger scale[37, 38], and methods like current-induced annealing are limited to electrically conductive materials and require electrode attachment[39, 40, 41]. Therefore, thermal solutions for sample cleaning are often used due to their rather simple installation and ease of use. Many setups allow annealing in controlled ambient environments, such as inert, reducing, and oxidizing atmospheres, as well as vacuum, the selection of which can have a drastic effect not only on the cleaning process but also on the integrity of the 2D material, with electrical doping[42, 43], defect formation[44, 45], induced strain[46, 47], and graphitization of carbonaceous contamination[48] being common side effects of thermal annealing processes.
@@
For graphene, even at temperatures as high as 500 °C, inert atmospheres have been reported to be ineffective at completely removing PMMA residue[49], and can turn it into covalently bonded amorphous carbon, which makes it difficult to remove[50, 30]. Reducing atmospheres can facilitate depolymerization of polymer residue when annealed at 400 °C or higher[42], with H2 and Ar/H2 being particularly effective in removing residual conjugated carbon systems by transforming them into light hydrocarbons[51]. However, this tends to induce rehybridization of graphene from sp2 to sp3, resulting in amorphous carbon[44, 46], as well as strong p-type doping[43]. Annealing in oxidizing atmospheres, while being effective in decomposing amorphous carbon contamination[52], can cause the formation of cracks in graphene at temperatures as low as 200 °C[44], and still fails to remove all contaminants, even at 500 °C[38]. Also, it has been found that the oxidative strength of the atmosphere has a strong effect on the etching characteristics, as NO2 seems to completely etch graphene away when annealed above 500 °C, while CO2 preserves it[45]. Thermal annealing in vacuum, while reported to be unable to remove contamination completely[2, 53, 27, 49], seems to introduce the least amount of lattice defects compared to other thermal annealing treatments due to the lack of reactive species in the annealing atmosphere[25]. In addition, the amount of induced hole doping and strain is reported to be minimal[43, 47], while providing satisfactory cleaning results[49, 38]. The purity of the vacuum must be kept in mind though, as annealing under better vacuum conditions reportedly yields cleaner specimens[50]. Pressure values in the range have been shown to be ineffective at reliably cleaning graphene surfaces[49], compared to annealing in high vacuum[25] or ultra-high vacuum (UHV)[2, 53, 27, 47], where large atomically clean surfaces have been achieved[38], making UHV annealing a suitable candidate for contamination removal without harming the integrity of the underlying material.
@@
Despite this established understanding of polymer degradation as well as hydrocarbon species desorption[54, 55] on graphene surfaces, there is still a lack of reporting on the cleaning of monolayer h-BN despite its various projected applications, such as being a dielectric for nanoelectronic devices[56, 57, 58, 59, 60], and for quantum information and sensor technology due to its quantum light emission across a wide spectral range[61, 62, 63, 64, 65]. Particularly, the latter relies on atomically clean surfaces to reliably conduct atomically-precise defect creation[66, 67, 68] or carbon doping[69, 70, 71] to create active sites for single photon emission. While surface contamination reduction via thermal annealing of h-BN has previously been utilized, including annealing in oxidizing environments[72, 73, 74, 75], and UHV [76, 77], the reports either lack information on the structure and abundance of contaminants as a function of annealing temperature, or do not employ atomically resolved characterization techniques, leaving open the possibility of a thin layer of contaminants remaining on the surface.
@@
In this work, we investigate the effect of thermal annealing in UHV on CVD-grown and wet-transfer-assisted free-standing monolayer graphene and h-BN, using a custom-built UHV heating chamber that is connected to a scanning transmission electron microscope through a vacuum transfer line. This ensures that the cleaning effect can be directly characterized without introducing airborne contaminants in the process. Contrary to commonly used spectroscopic assessment methods with limited spatial resolutions, including Raman spectroscopy[43, 30], X-ray photoelectron spectroscopy [2, 42, 49, 35], and time-of-flight secondary ion mass spectrometry [27, 47], we employ atomically resolved scanning transmission electron microscopy (STEM)to identify the presence of contamination on monolayer graphene and h-BN surfaces as a function of temperature down to the atomic level. Using annular dark field (ADF)-STEM imaging, we can identify the position, shape, and abundance of contamination due to its higher scattering intensity compared to atomically clean free-standing monolayer regions. Using contrast-based image thresholding and segmentation to separate contaminated from clean areas, we determine the relative amount of surface area covered by contamination, a procedure that is carried out for different transfer methods and baking temperatures, correlating the relative clean monolayer area of graphene and h-BN with annealing temperature. This is complemented with electron energy loss spectroscopy (EELS)measurements before and after the annealing steps, giving insight into the composition of the contamination that survives the heating process, which is one of the main limitations for obtaining truly residue-free materials.
@@ Methods
After investigating the sample quality and observing a large amount of metallic residue on all samples transferred this way, both before and after the annealing steps, further graphene samples were prepared via electrochemical delamination[78] using commercial monolayer graphene on Cu with PMMA coating (Graphenea, Inc.). To delaminate the samples, after cutting them to size as described above, a platinum anode was placed into NaOH (SigmaAldrich, ACS reagent, , pellets) aqueous solution (1 m), with the Cu/graphene/PMMA stack on Inox 02 tweezers acting as the cathode. All samples were delaminated at a voltage of 4 V. Afterwards, the delaminated graphene/PMMA stack was washed in DI-water to remove residual NaOH contamination, after which it was fished out with a TEM grid with a Quantifoil holey amorphous carbon membrane (Quantifoil R 1.2/1.3 Au grid), followed by a hot acetone bath for 1 h, with a final wash in IPA for 45 min. One additional sample was transferred onto a holey silicon nitride support film (PELCO, diameter, silicon support structure, 0.5×0.5 mm window, 200 nm membrane thickness, pore size) using the same transferring steps.
Hexagonal boron nitride— After sourcing monolayer h-BN from three different suppliers, we chose monolayer h-BN on copper from SigmaAldrich due to their superior monolayer coverage (Figure 15, Supporting Information), cut into small squares, which was spin-coated with PMMA (ALLRESIST AR-P 642.04, 200K) for 15 s at 500 RPM, followed by 15 s at 1000 RPM, and finally 60 s at 2000 RPM, at a ramp of 100 RPM/s. The Cu/h-BN/PMMA stack was placed on a glass slide, where the PMMA was left to cure on a hot plate at 150 °C for 30 min in air. After the PMMA had hardened, less than 1 mm was cut from all edges of the square to remove PMMA spillover to the other side of the foil. The subsequent delamination and washing steps were analogous to the graphene samples transferred via electrochemical delamination as described above.
@@
All annealing experiments were carried out in a custom-built UHV heating chamber integrated into the CANVAS system at the University of Vienna[97], with a base pressure typically around or lower. The samples are annealed via resistive heating of a tungsten filament that is placed out of the line of sight of the samples to avoid locally high temperatures and deposition of material coming off the wire. A more thorough description, including design sketches and used materials, can be found in the Supporting Information.
@@
We use commercial CVD-grown monolayer graphene and h-BN, which were transferred onto TEM grids using a PMMA-assisted wet transfer process (see Methods). The majority of the PMMA residue was removed from the surface by placing the samples into an acetone bath, followed by a further washing step in IPA. After preparation, the samples were inserted into the CANVAS system at the University of Vienna[97], which allows transporting the samples between a customized aberration-corrected scanning transmission electron microscope (Nion UltraSTEM 100)[80], and a custom-built UHV heating chamber without breaking the near-UHV conditions in between. This is integral to assessing the samples’ cleanliness since this allows the cleaning effect to be directly characterized without intermediary air exposure. Upon insertion into the loadlock of the CANVAS system, they were subjected to a routine bake at 160 °C for over 10 h in high vacuum to remove water and reduce the amount of surface contamination to avoid contaminating the UHV system. Using the Nion UltraSTEM 100, all samples were investigated for coverage and cleanliness before the UHV annealing step, after which they were transferred to the UHV heating chamber and annealed at different temperatures for 3 h. After the baking procedure had finished, the samples were immediately returned to the microscope for post-heating characterization to minimize the effect of electron-beam-induced hydrocarbon deposition[81].
To assess the cleanliness of the samples, both before and after the thermal treatment, up to ten typical holes ( nominal diameter) in the sample support were imaged. The holes were chosen based on average and characteristic features resulting from each heating process. Different holes in the sample support were selected before and after each heating step, as the dissociation and subsequent cross-linking of the hydrocarbon contaminants from prior e-beam exposure interferes with heating-induced contamination removal[82]. Approximately 20 arbitrarily chosen, non-overlapping areas with nominal FOV of 128×128 nm2 were captured within each hole, acquiring a total of ca. 150 - 200 frames per temperature and sample. At each new frame position, the electron beam was manually re-focused onto the sample surface to obtain lattice resolution, minimizing out-of-focus features in the contamination at the interface to the material’s surface. We note that no electron-beam-induced hydrocarbon deposition was observed during any measurement, and refer to the Supporting Information for a more thorough discussion.
@@
All as-prepared Easy Transfer Graphene samples exhibited a high amount of metallic contamination in the form of nanosized iron and copper clusters in addition to carbonaceous contamination covering practically the entire surface. While atomically clean monolayer regions could be found, they were small in size, typically less than 10 nm in diameter, and accounted for less than 5% of the free-standing regions (Figure 1a, blue dots). Example areas of contaminated and cleaned Easy Transfer Graphene are shown in Figure 1b. Heating to 250 °C did result in some removal of organic residue, increasing the average exposed clean area to 35%. The sizes of the now uncovered clean areas remained small (around 20 - 30 nm in diameter), indicating that higher temperatures are needed. Heating to 320 °C uncovered more of the surface area (ca. 40%), however, we identified thin hydrocarbon contamination networks all over the surface with only a marginal increase in the uncovered clean area compared to the 250 °C treatment. While annealing at 450 °C removed the majority of the carbonaceous contamination, resulting in a 70% relative clean monolayer area, with clean regions wider than being common, we identified networks of metal clusters embedded in carbon-rich contamination as the primary residual contamination. Raising the annealing temperature to 550 °C did not increase the unveiled clean area, which remained at around 70%. This not only indicates that the achievable cleanliness via UHV annealing is heavily limited by the residual metal contaminants introduced during the sample growth and transfer process, but also that temperatures above 400 °C appear to be sufficient to remove almost all of the carbonaceous contamination that is not directly bound to metal clusters.
@@
The remaining residues after annealing at 450 °C consist mainly of copper, but also iron nanoclusters (Supplementary Figure 12, Supporting Information), embedded in carbon-rich contamination, primarily located at graphene grain boundaries, shown in Figure 2a. These metal clusters also form aggregates with a higher density of metallic residue, which are not necessarily attached to grain boundaries as seen in Figure 2b, from which dendritic carbon-rich structures appear to emerge. Supplementary Figure 13 (Supporting Information) shows that the metal contaminants are rich in Fe and Cu. We also identify nanosized holes in the material with typical sizes between 15 and 20 nm (Figure 2c). However, based on the relatively low abundance of these holes, we presume that they were already present in the material before annealing and were covered by the contamination that was removed during the heating process, instead of being annealing-induced damage. Atomically clean areas that would otherwise be thousands of nm2 in size contain some isolated small carbon contamination patches (Figure 2d), likely attached to point defects in the material. Contaminants in regions where graphene layers folded over on themselves during the sample transfer (Figure 2e) are difficult to remove due to a rise in surface energy with increasing layer number[83], which inhibits surface contamination desorption. Additionally, contaminants could not only be on the surfaces but also in between the layers, which would make them difficult to remove, as they cannot penetrate the layers they are sandwiched between. This results in typically dirtier areas compared to monolayer regions.
@@
Upon investigating the pre-cleaned samples, we observed that the quality of the transferred h-BN was inferior to that of graphene made via the same transfer method. Firstly, the suspended regions showed more holes and folds compared to graphene, likely originating from the transfer process. Secondly, a large fraction of the surface consisted of multilayer structures, which appeared to emerge from SiO2 nanoclusters (Supplementary Figure 14, Supporting Information). We presume that they were introduced during the sample growth, possibly from evaporation of SiO2 from the quartz tube during the CVD growth process[84], rather than during storage or transfer. Together, these two sources contributed to an estimated monolayer coverage of only around 50% (Supplementary Figure 14, Supporting Information), the highest among the three h-BN suppliers considered for this study (Supplementary Figure 15, Supporting Information), compared to practically 100% for graphene, despite the similar sample preparation procedures. This is also reflected in the strong variation of measured clean monolayer area (Figure 3a) across all temperatures, which is largely due to differences in monolayer coverage per imaged area, rather than local variations in residual contamination at a given temperature.